### v2

# I. Márquez 11    C. Boisson 22    M. Joly 22    D. Pelat 22    F. Durret 33

Abstract
Context. The relationship between an active galactic nucleus (AGN) and its host galaxy is still far from being understood. Properties of the host galaxies of Seyfert nuclei, such as luminosity concentration, morphological type, metallicity, and age of the stellar population, are expected to be related with nuclear activity – either at the epoch of galaxy formation or at the present day via feeding of the central black hole.
Aims. In this paper, we investigate whether stellar ages and metallicities are linked to the activity within the nucleus in a sample of AGNs of various types.
Methods. Our sample includes seven AGNs, from Seyfert 1 to LINERs, observed with VLT/ISAAC and VLT/SINFONI. Based on an inverse method using a stellar library, we analysed H band infrared spectra, in a wavelength region devoid of emission lines, at a spectral resolution of $\approx$3000, in the few central 100 pc. Hubble Space Telescope images were used to visualise the regions defined in each galaxy.
Results. For each galaxy, we give the results of the spectral synthesis, in particular the percentages of the stellar, power law, and black-body continua, and the percentages of various stellar types that account for the stellar lines.
Conclusions. Out of the seven galaxies, three show strong and recent star formation in the inner 100 pc, while no star formation is detected in the three genuine Seyfert 2 galaxies. Beyond a radius of 100 pc, all show more or less recent star formation. Moreover, we can conclude that the star formation history of the inner nucleus is highly heterogeneous.
Key Words.:
Galaxies: active.
1 Introduction
An important yet unsolved issue for understanding the Seyfert phenomenon is the relation between the host galaxy and the properties of activity within the nucleus. It is now accepted that nuclear star formation and central engine activity can coexist in the inner regions of the active galactic nuclei (AGNs) host galaxy, as derived from the strong relation linking the mass of super-massive black holes with the velocity dispersion of their host galaxy bulge (e.g. Gebhardt et al., 2000; Kormendy & Ho, 2013). The accelerated star formation caused by dynamic instabilities that trigger and/or fuel nuclear activity could result in an overabundance of giant, supergiant, and supermetal-rich (SMR) stars (e.g. Scoville et al., 2023, and references therein).
Stellar population syntheses of the cores of AGNs provide information on stellar masses, mean stellar ages, abundances, and dust content, as well as insight on their star-formation history. Such information is derived from spatially unresolved galaxy spectra and thus reflect global averages over properties that are known to vary across the galaxies, but still provide valuable insights. Observations of the cores of nearby active galaxies allow the complex star formation history in active galaxies to be exhibited (e.g. Terlevich et al., 1990; Storchi-Bergmann et al., 2001; Diniz et al., 2017; Mallmann et al., 2018; Dametto et al., 2019). Studies in the optical and infrared (IR) domains have indicated that young or intermediate age populations can contribute significantly to AGN central spectral emission (e.g. Oliva et al., 1999; Boisson et al., 2000; Chapman et al., 2000; Boisson et al., 2004; García-Bernete et al., 2015; Thomas et al., 2018; Burtscher et al., 2021; Dahmer-Hahn et al., 2022; Riffel et al., 2024, and references therein).
To study the properties of the stellar population in the inner few tens of parsecs, all studies face the problem of disentangling the stellar population from the other components: the continuum emission by the AGN, the emission from heated dust, and galactic extinction. Contrary to the optical spectral range, near-IR allows one to access to regions highly obscured by dust, and exhibits many stellar features that permit one to constrain the average spectral type, luminosity class, and metallicity markers characteristic of a large variety of stellar populations (e.g. Origlia et al., 1993; Dallier et al., 1996; Ivanov et al., 2004; Stephens & Frogel, 2004). Since part of the near-IR H-band is free of the strongest emission lines due to the AGN, it is a well-suited domain for the determination of the stellar content of central regions of Seyfert galaxies, including Seyfert 1s, as long as the spectral resolution is high enough (e.g. Dallier et al., 1996; Frémaux et al., 2006; Cesetti et al., 2009).
[…]
In Section 2 of the paper, we briefly present the sample of seven nearby AGNs observed with the VLT and summarise the data reduction. The spectral population synthesis code and its implementation are presented in Section 3. Results for individual AGN are presented in Section 4 and discussed in a global context in Section 5.
2 The data
To study the differential stellar population in the cores of AGNs, we assembled a sample of seven sources typical of their class of activity, three Seyfert 1 galaxies (MCG-6-30-15, NGC 2992 and NGC 3783) and four Seyfert 2s (NGC 2110, NGC 3185, NGC 6221 and NGC 7582). Five galaxies observed with the VLT/ISAAC spectrograph, were previously presented in Boisson et al. (2002) and are complemented by two more galaxies (this paper) observed with the same equipment. We combined these data with VLT/SINFONI spectra kindly provided by Sandra Raimundo for MCG-6-30-15 (Raimundo et al., 2013).
Hubble Space Telescope (HST) archival images were used to visualise the different regions as selected for each galaxy. Informations on the seven galaxies of this study are given in Table 1. In this paper we assume H0 = 70 km s-1 Mpc-1, $\Omega_{m}$=0.3, $\Omega_{\Lambda}$=0.7.
Table 1: Galaxies observed with Isaac.
Name Sy Morphology RA DEC z E(B-V) PA HST D scale
type (J2000.0) (J2000.0) ($\deg$) filter (Mpc) (pc/”)
MCG-6-30-15 1 E/S0 13h35m53.7s -34d17m44s 0.00775 0.06 86 F547M 31 150
NGC 2110 2 E3/S0 05h52m11.4s -07d27m22s 0.00779 0.37 $-25$ fr680p15 32 157
NGC 2992 1/2 Sa 09h45m42.0s -14d19m35s 0.00771 0.06 55 F196N 31 150
NGC 3185 2 SBa 10h17m38.5s +21d41m18s 0.00406 0.03 0 F814W 16 78
NGC 3783 1 SBa 11h39m01.7s -37d44m19s 0.00973 0.12 $-77$ F160W 40 189
NGC 6221 2 SBbc 16h52m46.1s -59d13m07s 0.00500 0.17 30 F160W 19 92
NGC 7582 2 SBab 23h18m23.5s -42d22m14s 0.00525 0.01 0 F606W 22 107
The columns are: 1) name, 2) Seyfert type, 3) morphological type, 4) and 5) coordinates, 6) redshift, 7) Galactic E(B-V), 8) slit position angle, 9) HST filter, 10) distance, 11) scale in parsecs per arcsecond. Galactic E(B-V) values are from Schlafly & Finkbeiner (2011) as given by the Dust Extinction tool provided by IRSA (https://irsa.ipac.caltech.edu/applications/DUST/).
2.1 VLT/ISAAC data
[…]
The slits were always centred on the IR galaxy nucleus. Their position angles are provided in Table 1.
The data reduction is fully described in Boisson et al. (2002), including correction for electrical ghosts, flat-field, distortions along and perpendicular to the slit, and wavelength calibration. Early type stellar templates were used for the correction of telluric absorption. The spectra were extracted by using logarithmic flux profiles to define a nucleus and regions around it. All of the spectra displayed are shifted to the rest wavelength. Whenever spectra on each side of the nucleus were very similar, a pseudo ring spectrum was created by summing both individual spectra, allowing for a better signal to noise ratio.
2.2 VLT/SINFONI data
[…]
2.3 HST images
Archival WFPC2 HST images are available for all the galaxies (see the HST filters in Table 1). The reddest high quality images were used. In order to enhance morphological features in the innermost few arcseconds of the galaxies, we created sharp-divided images by dividing the original images by their smoothed counterparts (Márquez et al., 2003). This enhanced any structural inhomogeneities that could affect the stellar populations, such as a dust component.
The spectral regions extracted from ISAAC and SINFONI observations were drawn on the images, and the slit width is 1″. North is up and east to the left.
3 Stellar population synthesis
The analysis of the stellar populations inside unresolved cores of active galaxies has been the aim of several studies for some time now. All studies face the problem of disentangling the stellar population from all the other components: the continuum emission by the AGN, the emission from the heated dust, and galactic extinction.
The near-IR range is a very useful domain to study the stellar content of active galaxies. In particular, the H-band is well suited, as the non-stellar contribution (mainly reprocessed nuclear emission by dust) is much less than in the K-band window. The H-band is also the region in which emission from cool stars peaks, and the spectrum is rich in luminosity sensitive stellar absorption lines, leading to the age of the stellar population (e.g. Dallier et al., 1996; Frémaux et al., 2006). Extinction by dust is lower there than for the shorter wavelength J-band.
Since part of the IR H-band is free of the strongest AGN emission lines, it may allow for the determination of stellar populations in the very internal regions of Seyfert galaxies in the presence of a diluting continuum, as is the case for Seyfert 1s. In particular the spectral range considered in this paper, covering 1.57 to 1.64 $\mu$m, includes many metallic stellar features (FeI, MgI, SiI, CaI, NiI, CO, and OH). The spectral variations of the stellar component are generally complex, and the age and the metallicity of the stars may have similar influences on the integrated spectrum. Therefore, it is crucial to use as many observables as possible to decrease the degeneracy. Indeed the near-IR CaII triplet and the IR CO and Si indices can be strong not only in supergiant stars but also in metal-rich giants; colour gradients can be caused either by changing stellar populations or by dust effects.
There are essentially two different approaches to differential population synthesis. The ’direct’ approach, also called the ’evolutive synthesis method’ (e.g. Tinsley, 1972; Bruzual & Charlot, 2003; Le Borgne et al., 2004; Vázquez & Leitherer, 2005), compares the spectrum at each time step of an evolving stellar population with the spectrum of the studied galaxy. The other method aims to extract the stellar components from the spectrum of the astrophysical object through a mathematical inversion of the information content of the observed spectrum. This so-called inverse method uses building blocks of either stellar spectra (e.g. Faber, 1972; Pickles, 1985; Pelat, 1997, 1998; Moultaka & Pelat, 2000; Frémaux et al., 2007) or simple stellar populations (SSPs; e.g. Bica, 1988; Vergely et al., 2002; Cid Fernandes et al., 2005; Gomes & Papaderos, 2017; Wilkinson et al., 2017; Cid Fernandes, 2018). In this paper, we use an inverse method, based on the minimisation of a quadratic form subject to constraints in the form of inequalities and equalities, developed and thoroughly tested by Frémaux et al. (2007).
Taking the flux received in each spectral pixel as the observable quantity to fit increases the number of observables, allowing for a greater sampling for the stellar composition of the studied galaxies, and alleviating the difficulties to determine the level and shape of the continuum. The specificity of the method is to take the non stellar parameters (reddening, dust emission, and AGN non-stellar continuum) directly into account in the synthetic distance to be minimised. These non-stellar parameters introduce non-linearity into the objective function, expressed as a quadratic distance to minimise. The lower this value, the closer from the observation the synthetic spectrum.
Due to degeneracy of these components, mainly in low S/N ($<$ 40), the non-stellar parameters are bound to values commonly deduced from astrophysical considerations. Dust grains with a temperature below 200 K mainly contribute to the flux spectrum farther in the IR, so their contribution in our wavelength range is too small to be detected. A temperature of 1500 K is the maximum possible for dust grains as it corresponds to the sublimation temperature of the most refractory dust material. Most of the active galaxies have an intrinsic reddening E(B-V) smaller than 3. For each galaxy the minimum allowed galactic extinction was fixed to the Galactic value on the line of sight (Schlafly & Finkbeiner, 2011). The AGN continuum power law index $\alpha$ is the most difficult non stellar parameter to estimate. Indeed, in our very small wavelength domain, variations of $\alpha$ only induce a very weak variation in the slope of the fitted continuum, as explained in Frémaux et al. (2007). We therefore fixed $\alpha$ to the typical value of 1.5 for AGNs222the power law is defined by $\nu^{-\alpha}$ (Vanden Berk et al., 2001). However, these bounds can easily be modified for the synthesis of individual objects.
Such a method is inherently prone to errors due to the presence of noise in the data and potential degeneracies inherent to the studied problem. To assess the impact of these errors on the estimated parameters, we performed a series of Monte Carlo simulations. Gaussian random noise with zero mean and a standard deviation of 10% of the observed galactic spectrum value was added to each pixel of the spectrum. This corresponds approximately to the signal-to-noise ratio of the data. For each object, 300 independent Monte Carlo trials were conducted. To mitigate the influence of outliers, 38% of the highest and 38% of the lowest values for each parameter were excluded. This percentage corresponds to the optimal truncated estimator for the worst-case scenario, that is noise with an undefined mean (Cauchy noise). The standard deviation of the remaining values provides the error bars for the parameters obtained via the minimisation method.
Whatever the approach in stellar syntheses, a spectral database is needed to build a synthetic galactic spectrum. The stellar libraries come in two flavours, either built with theoretical or empirical spectra. Though theoretical libraries could allow for a wide range and combination of stellar parameters, it requires exhaustive model atmospheres that depend on atomic and molecular line opacity data bases not always available.
As demonstrated by Frémaux et al. (2006), the comparison between observed and computed spectra deteriorates for decreasing temperature, where several molecular as well as metallic absorption lines are still missing in atmospheric models. Working in the IR H-band, we thus use an empirical stellar library. It should be noted that each individual spectral type in a library encompasses a significantly large domain of temperature, gravity, and metallicity, because one cannot have as many types of stars in the database as those forming a galaxy.
[…]
Also, the finite spectral resolution of observations imposes a limitation in the extent of the library due to possible degeneracy among stellar spectra. To avoid multiple degenerate solutions for the minimisation problem, the library is composed of 30 stars of all luminosity classes and of spectral types from A0 to M4, chosen to describe at best the HR diagram (see discussion in Frémaux et al., 2007). The code includes physical constraints on the relative number of born stars of the different mass groups, here in the case of a standard Initial Mass Function (see a discussion in Moultaka et al., 2004). Table 2 gives the spectral type, absolute visible magnitude, temperature, luminosity, mass, and age of each of the stellar types considered together with the name of a typical star. When known, the metallicity ([Fe/H]) of this star is provided with the associated reference. Although we do not intend to give the exact metallicity of each stellar component of a galaxy, we should be able to distinguish qualitatively a metal-rich from a solar metallicity population. In the following we refer to super metal-rich stars as SMR.
In the text, we define as young, intermediate and old stellar populations those with stars with ages $\leq$ 107 yrs (supergiant stars), around 108 yrs (A0-F2V), and older than 109 yrs (dwarfs and giants) respectively. The supergiant stars are indeed used as a proxy for a young population since no stars hotter than A type can be considered; the contribution of O and B type stars is negligible in the IR, compared to cold or intermediate stars (A, F spectral types). For more details on the spectral fitting code and database see Frémaux et al. (2007).
Table 2: Stellar parameters.
Name Spectral MV T(K) L/L☉ M(M☉) Lifetime [Fe/H] Ref.
Type (yrs)
HD172167 A0V 0.5 9500 54 3.0 2.7$\times$108
HD48501 F2V 3.6 7100 2.9 1.5 3.9$\times$108 0.01 BT
HD112164 rG1V 4.5 5900 1.0 1.0 6.0$\times$109 0.24 E
HD10307 G1.5V $>$4 5880 1.0 1.0 6.2$\times$109 0.0 L
HD106116 rG4V 4.9 5700 1.0 1.0 6.9$\times$109 0.15 An
HD185144 wK0V 6.0 5300 0.42 0.8 6.4$\times$109 -0.18 R
HD39715 K3V 6.65 4740 0.32 0.7 1.1$\times$1010 -0.04 VF
HD131977 K4V 7.2 4600 0.2 0.66 3.5$\times$1010 0.05 L
HD201091 wK5V 6.0 4400 0.32 0.7 3.2$\times$1010 -0.35 KV
HD79210 M0V 9.0 4000 0.09 0.5 $>$3$\times$1010 0.0 M
HD36395 rM1V 9.3 3700 0.07 0.45 $>$3$\times$1010 0.49 M
HD119850 wM1.5V 9.3 3700 0.07 0.45 $>$3$\times$1010 -0.3 M
HD173739 wM3V 11.5 3700 0.02 0.25 $>$3$\times$1010 -0.29 M
HD27397 F0IV 2.7 7000 6.5 1.3 2.1$\times$109
HD115604 rF3III 1.6 6800 17 1.5-1.6 3.0$\times$109 0.18 H
HD220657 F8III 1.2 5800 30 1.2 6.0$\times$109
HD107950 G6III 0.85 5000 40 1.0 9.3$\times$109 0.08 ASF
HD206952 rK1III 0.6 4700 70 1.1 6.0$\times$109 0.2 P
HD3627 K3III 0.5 4300 110 1.10 6.0$\times$109 0.04 MLSF
HD3346 K5-M0III -0.2 3590 330 1.1 6.0$\times$109
HD102212 wM1IIIab -0.5 3800 430 1.25 6.0$\times$109 -0.3 S
HD151732 M4III -0.5 3535 880 1.3 4.0$\times$109
HD19058 wM4II -1.2 3500 450 5 1.1$\times$108 -0.15 SL
HD36673 F0Ib -6.6 7400 30000 12 1.8$\times$107 0.04 V
HD20902 F5Ib -6.6 7000 30000 12 1.8$\times$107 -0.05 P
HD139717 rF8Ib -6.5 6500 30000 9 2.5$\times$107 0.15 Ad
HD204867 rG0Ib -6.4 5500 30000 10 1.4$\times$107 0.10 KV
HD210745 rK2I -5.95 4300 30000 12 1.7$\times$107 0.17 KV
HD201251 rK4I -5.8 4000 33000 12 1.8$\times$107 0.12 Lu
HD39801 M2I -6.0 3500 63000 18 8.5$\times$106 0.01 P
r subscript for SMR; w subscript for below solar metallicity
References: An: Andrievsky et al. (2013); Ad: Adibekyan et al. (2012); ASF: Afşar et al. (2012); BT: Boesgaard & Tripicco (1987); E: Edvardsson et al. (1993); H: Hauck et al. (1985); KV: Koleva & Vazdekis (2012); L: Lee et al. (2011); Lu: Luck (2014); M: Mann et al. (2015); MLSF: Massarotti et al. (2008); P: Prugniel et al. (2011); R: Ramírez et al. (2013); S: Sheffield et al. (2012); SL: Smith & Lambert (1986); V: Venn (1995); VF: Valenti & Fischer (2005).
4 Spectral fits
Synthetic stellar populations are computed for all the individual spectra extracted from the galaxy frames, using the stellar database previously described.
The velocity dispersions, $\sigma$, of the galaxies observed with ISAAC, were measured by cross-correlation of the full observed H-band spectrum of the galaxy with IR template stars (see Boisson et al., 2000, 2002). The resulting $\sigma$ values are given in Tables 3 to 9. The differential velocity dispersion between galaxies and template stellar spectra is handled in the synthesis.
The synthesis results are given in Tables 3 to 9, and are displayed in accompanying figures showing the observed and computed spectra together with the fractional difference between the model and data. The observed spectra are shown in black, normalised by the flux around 1.586 $\mu$m, with the synthesis shown in red. The radius of the nuclear region sampled, and the width of the off-nuclear zones are given in parsecs. The various parameters listed are: first, the percentage of stellar (stell), power law (PL) and black-body (BB) emissions, together with the corresponding black-body temperature T(BB), and the intrinsic reddening E(B-V); second, the stellar composition of the galaxy, together with the percentage of contribution of each stellar type to the galactic spectrum at the reference wavelength. To ensure readability only stars having a contribution larger than 1% are listed.
4.1 MCG-6-30-15
Table 3: Spectral synthesis of MCG-6-30-15.
SINFONI SINFONI SINFONI SINFONI SINFONI ISAAC ISAAC ISAAC
E(1) N(2) W(3) S(4) ring C(nuc) E W
60-100 pc 20-60 pc 60-100 pc 20-60 pc 20-170 pc 110 pc 110-330 pc 110-330 pc
$\sigma$ km/s 80 80 80 80 80 70 80 80
% stell 41$\pm$1 23$\pm$1 47$\pm$1 19$\pm$1 39$\pm$1 17$\pm$1 100$\pm$0 100$\pm$0
% PL 57$\pm$2 74$\pm$1 52$\pm$1 76$\pm$1 57$\pm$1 83$\pm$1 0 0
% BB 2±$\pm 1$ 3$\pm$0 1$\pm$1 5$\pm$1 4$\pm$0 - - -
T (BB) 450$\pm$190 250$\pm$10 210$\pm$13 330$\pm$20 290$\pm$6 - - -
E(B-V) 1.08$\pm$0.2 - 1.27$\pm$0.01 - - 0.47$\pm$0.04 - 1.6$\pm$0.1
F2V 16$\pm$3 8$\pm$3 49$\pm$1 23$\pm$1
rG1V 14$\pm$10 13$\pm$8 10$\pm$3
rG4V 22$\pm$4 9$\pm$10
K4V 6$\pm$2 2$\pm$7 11$\pm$1 11$\pm$3 14$\pm$3
M3V 17$\pm$1 28$\pm$1
F8III 5$\pm$1
K5-M0III 29$\pm$1 49$\pm$1
M1IIIab 4$\pm$13 4$\pm$3
M4III 75$\pm$1 40$\pm$2 37$\pm$1 44$\pm$4 36$\pm$17 68$\pm$3
wM4II 2$\pm$1
rF8Ib 8$\pm$1 9$\pm$2 7$\pm$1
rK4I 2$\pm$1
M2I 6$\pm$1 18$\pm$2 27$\pm$1 32$\pm$4 24$\pm$2 18$\pm$3 3$\pm$1
The numbers given for the main contributing stars are the percentages of the various contributions to the galaxy spectrum at the reference wavelength 1.586 $\mu$m, as in all following tables.
Figure 1: Sharp-divided optical HST image in the F547M filter of MCG-6-30-15. The region studied by Raimundo et al. (2013) that we call ’pseudo-ring’ corresponds to the large blue circle (1.1 arcsec radius) minus the small blue rectangle, and their regions labelled 1 to 4 are shown in red. The ISAAC slit is superimposed in green, with the centre and the two regions shown as the central and external rectangles respectively.
Figure 2: Stellar population fits to regions 1-4 of MCG-6-30-15 (see Fig. 1 top).
Figure 3: Stellar population fit to the ’pseudo-ring’ of MCG-6-30-15 (see Fig. 1 top).
Figure 4: Stellar population fits to the three regions of MCG-6-30-15 taken with VLT/ISAAC. From top to bottom: nucleus, east, and west (see Fig. 1 bottom)
The Seyfert 1 galaxy MCG-6-30-15 has been extensively studied in all wavelength bands though primarily in X-rays. The host galaxy is an elongated E/S0. Photometric colour maps show a dust lane crossing the galaxy just south of the nucleus, roughly parallel to the major axis of the galaxy (Ferruit et al., 2000).
From a spectral analysis in the optical, the bulge ($\sim$ 200 - 800 pc) stellar population has been found to be dominated by old mildly metal-rich stars with a contribution of intermediate age stars (around 1 Gyr) that increases farther away from the galactic centre, see Boisson et al. (2004).
With its high spatial resolution, SINFONI allows one to sample the inner nuclear regions well. We analyse the five SINFONI regions (1 to 4 and a pseudo-ring of radius $\sim$1.1 arcsec excluding the very nucleus) observed by Raimundo et al. (2013). The ISAAC spectra sample two more regions besides the ones mapped by SINFONI: the nucleus ($<$ 1 arcsec /110 pc radius) and two regions east and west. The regions studied are marked on the sharp-divided image of the central region of MCG-6-30-15 as shown in Fig. 1.
The results of the stellar population fits of the various regions of MCG-6-30-15 are given in Table 3, with the corresponding spectra and fits shown in Figs. 2 and 3. We note the excellent agreement between the observed and synthetic spectra.
The presence of a non-stellar component is important in all SINFONI regions, specially in regions 2 (N) and 4 (S), which are closer to the nucleus, and therefore may be more contaminated by the AGN emission. Contrary to the SINFONI spectra, the inner 110 pc ISAAC spectrum includes the very nucleus, which then induces a higher contribution of the power-law component (83%).
The whole central $<$110 pc radius population is old ($\sim$ 5.0 $\times$109 yrs), dominated by dwarfs and evolved M giant stars. An important contribution of young (107 yrs) supergiant stars (M2I) is also present. Using emission lines to study the supernova rate, Raimundo et al. (2017) also deduce the presence of a young $<$108 yrs population. We note that the better spectral resolution of SINFONI, compared to ISAAC, reveals the presence of old super metal-rich dwarf stars as expected in case of successive bursts of star formation.
As can be seen from the ISAAC data, the outer regions (110-330 pc) are dominated by intermediate (F2V, $\sim$ 4.0 $\times$108 yrs) and old stars with no recent burst of star formation as indeed expected in lenticular galaxies (see Fig. 4). At such a large distance from the nucleus, the metallicity of the contributing stars is weaker than in the nuclear regions. A $\sim$ 20% contribution of young star population is present in the nucleus ( $<$ 110 pc).
Analysing in a large IR range (J, H, and K bands) the inner 300 pc, Riffel et al. (2024) found a dominating old population, the youngest contribution they obtain is of intermediate age ($>$108 yrs). This may be due to a lack of spatial resolution.
In conclusion, the nuclear stellar content of the Seyfert 1 MCG–6-30-15 is dominated by a mixture of old and young populations with the contribution of metal-rich stars proving a history of multiple bursts of star formation. This is in contrast with the overall bulge population inferred from our study in the optical, and points to a population gradient with ongoing star formation in the nucleus.
4.2 NGC 2110
Table 4: Spectral synthesis of NGC 2110.
C(nuc) SE NW1 NW2
25 pc 25-120 pc 25-120 pc 120-235 pc
$\sigma$ km/s 100 100 100 100
% stell 50$\pm$1 100$\pm$0 100$\pm$0 100$\pm$0
% PL 50$\pm$1 0 0 0
% BB - - - -
T (BB) - - - -
E(B-V) 0.94$\pm$0.20 1.23$\pm$0.17 - -
A0V 3$\pm$1
F2V 26$\pm$1
rG1V 24$\pm$1 27$\pm$1 23$\pm$1
wM3V 28$\pm$4 13$\pm$1
F0IV 12$\pm$1 13$\pm$1 1$\pm$1
F8III 26$\pm$2 10$\pm$1 19$\pm$1
G6III 22$\pm$3
K5-M0III 2$\pm$1 33$\pm$1
M4III 52$\pm$1 24$\pm$4 28$\pm$1 14$\pm$1
Figure 5: Sharp-divided IR HST image in the fr680p15 filter of NGC 2110 with the ISAAC slit superimposed and the studied regions marked.
Figure 6: Stellar population fits to the four regions of NGC 2110. From top to bottom: the nucleus, the SE, and the two NW regions (see Fig. 5)
NGC 2110 is a weakly barred S0 galaxy harbouring a Seyfert 2 nucleus, also classified as a FRI radio galaxy (Beckmann & Do Cao, 2010). High spatial resolution HST continuum images show a dusty circumnuclear disk with a spiral pattern (Evans et al., 2006; Schnorr-Müller et al., 2014, see also Fig. 5).
The sharp-divided image of NGC 2110 is shown in Fig. 5 with the studied regions superimposed. We extracted the spectra of four regions: a nuclear region 50x150 pc2 (C), two regions on each side of the nucleus extending from 25 to 120 pc (SE and NW1) and a second adjacent region NW up to 235 pc from the nucleus (NW2).
The stellar population in the nucleus of NGC 2110 is dominated by giant evolved stars. In the ring surrounding the nucleus, up to 120 pc, an old population is still dominating, but includes partly metal-rich and partly weak metallicity dwarf stars. The farthest NW region, beyond 120 pc of the centre, shows a large contribution of intermediate stars indicative of an ancient ($\sim$ 3.0 $\times$108 yrs) star formation burst in the past history of the galaxy. This population gradient is in line with the results of a study of the inner 3x3 arcsec2 (450x450 pc2) by Diniz et al. (2019) through near IR J and K bands integral field spectroscopy.
We do not detect the strong signature of young stars in the central 0.6 arcsec as claimed from optical integral field spectroscopy by Dahmer-Hahn et al. (2022). Neither do Burtscher et al. (2021) in a similar study, who point to a largely dominating old stellar population. In the near IR, Riffel et al. (2024) found a stellar population of mean age $\sim$ 5 Gyr in a region that includes our three innermost regions, in global agreement with us. And though we detect sub-solar metallicity dwarfs as in Dahmer-Hahn et al. (2022), we also highlight a contribution of metal-rich stars in this same region.
The synthetic spectra are displayed in Fig. 6. It is worth noticing that, in the four regions analysed, residuals are clearly marked around 1.59 $\mu$m and 1.62 $\mu$m, where one expects a combination of several molecular features (SiI, FeI, OH, CO, CaI). Such line strengths and ratios vary with both stellar temperature and gravity, implying that ultra-cool dwarfs are missing in the data base we use. NGC 2110 is the only galaxy in the small sample studied where such a lack appears, implying a peculiar stellar population for this object.
In conclusion, the stars that contribute the most to the stellar population in the inner 235 pc are old and cold, indicating the absence of star formation since $\sim$ 6.0 $\times$109 yrs, although a younger population is present in the outer regions. The presence of a metal-rich star contribution shows that in the past there have been many star formation episodes. This was suggested in the optical study (Boisson et al., 2000) where, in the whole bulge up to 800 pc, an old but metal-rich stellar population is the proof of multiple star formation episodes in the past.
4.3 NGC 2992
Table 5: Parameters of the spectral synthesis for NGC 2992.
C(nuc) SW NE
110 pc 110-330 pc 110-330 pc
$\sigma$ km/s 95 80 90
% stell 38$\pm$1 64$\pm$2 98$\pm$0
% PL 62$\pm$1 0 0
% BB - 36$\pm$2 2$\pm$0
T (BB) - 970$\pm$23 200$\pm$0
E(B-V) 0.70$\pm$0.07 - -
F2V 19$\pm$1
rG1V 14$\pm$5 36$\pm$2 13$\pm$1
K4V 5$\pm$2
wM3V 14$\pm$1
F0IV 3$\pm$1
F8III 12$\pm$1
M4III 81$\pm$1 34$\pm$1 42$\pm$0
M2I 27$\pm$0
Figure 7: Sharp-divided IR HST image in the F196N filter of NGC 2992 with the ISAAC slit superimposed and the three regions indicated: the nucleus (C), the north-east (NE), and south-west (SW) regions.
Figure 8: Stellar population fits to the three regions of NGC 2992. From top to bottom: nucleus, SW, and NE (see Fig. 7).
[…]
In a detailed IR adaptive optics study, Chapman et al. (2000) found a radial distribution of the CO index in the inner 3 arcsec ($\sim$ 450 pc), suggesting a population gradient within the core, with the stellar population at the very centre being older than that in the surrounding regions. Stellar population modelling of OASIS integral field spectroscopy by Stoklasová et al. (2009) revealed a contribution of young stars close to the dust lane, younger than in the rest of their field of view, with mostly old stars in the eastern part. From optical GMOS observations, Guolo-Pereira et al. (2021) find a mainly old metal-rich stellar population, with a noticeable contribution of young metal-poor stars in the inner 1.1 kpc of NGC 2992. Burtscher et al. (2021) from VLT/X-shooter spectra in an aperture of 150 parsec radius infer a dominating rather old reddened stellar population ($>10^{9}$ yrs). Another study by Dahmer-Hahn et al. (2022) with MUSE integral field spectroscopy of the 270 central parsecs, points to a mainly young, metal-rich, and highly reddened stellar population to the north–west, with the rest of their field of view (FoV) having an old, metal-poor, and less reddened stellar population. However, a study in the IR by Riffel et al. (2024), with SpeX attached to the IRTF, sampling the same dimensions, infers a largely dominating old population ($\sim$ 8 Gyr).
In Table 5 the results of the stellar population fits for the three regions of NGC 2992 we studied are given, with the corresponding spectra and fits shown in Fig. 8. The AGN non stellar contribution reaches about 60% in the nucleus. We find an evolved (M giant stars, $4\times 10^{9}$ yrs) stellar population of about solar metallicity, in a high extinction environment, with no hint of a young stellar population contribution. The stellar population is younger in the regions surrounding the nucleus, with not more than the foreground galactic extinction. An important star formation difference is visible between the SW and NE regions: presence of an important metal-rich component and recent ($\sim 10^{7}$ yrs) starburst in the SW region, the closest to the dust lane, and an older burst of star formation inducing an intermediate age stellar population in the NE region.
The hot dust temperature inferred by the analysis of the SW region is not expected. A test constraining the dust temperature to 500 K induces an increase of the computed extinction up to the nuclear one, a plausible result as this region is the closest to the dust lane. The global stellar population remains the same, pointing to some degeneracy between extinction and featureless continuum from dust emission. In our study of the 600 central parsecs, taking advantage of the H-band, we can emphasise a clear age and possible metallicity gradient, besides a high star formation in the SW region.
4.4 NGC 3185
Table 6: Parameters of the spectral synthesis for NGC 3185.
C (nuc) R (ring)
60 pc 60-175 pc
$\sigma$ km/s 75 80
% stell 100$\pm$0 100$\pm$0
% PL 0 0
% BB - -
T (BB) - -
E(B-V) - -
F2V 10$\pm$0
rG1V 14$\pm$1 28$\pm$0
K4V 11$\pm$1
F0IV 18$\pm$1 14$\pm$0
F8III 4$\pm$1
M4III 53$\pm$1 48$\pm$0
Figure 9: Sharp-divided HST image in the F814W filter of NGC 3185 with the ISAAC slit superimposed. The nuclear region is indicated by C, and the ring corresponds to the sum of the two blue regions.
Figure 10: Stellar population fits to the two regions of NGC 3185. From top to bottom: nucleus and ring (see Fig. 9).
NGC 3185 is a strongly barred SBa galaxy located in a compact Hickson group dominated by NGC 3190. This galaxy classified as Seyfert 2 by Filippenko & Sargent (1985), is also considered as a transition object with both Seyfert and star forming characteristics (e.g. Gonçalves et al., 1999). A LINER classification was also proposed by Carrillo et al. (1999). Circumgalactic rings of star forming material are observed in both UV and IR images on a large scale (Lanz et al., 2013).
From optical studies, NGC 3185 has been shown to have a very old population in the central 1.56 kpc (James & Percival, 2016). To our knowledge, there are no other published results on the stellar content of the nucleus of this galaxy.
The sharp-divided image of the galaxy is shown in Fig. 9 with the studied regions superimposed. The spectra of the two regions on either side of the nucleus are very similar and thus combined to represent the spectrum of a ring-like feature at about 120 pc from the nucleus. The stellar population derived for the nucleus and ring are given in Table 6, with the corresponding spectra shown in Fig. 10.
Contrary to what is expected for a genuine Seyfert nucleus, no dilution by a power law component is needed to account for the nuclear spectrum. Within the central 360 pc of NGC 3185, the stellar population is old, partly metal-rich, though intermediate-age dwarf F stars in the ring indicate some stellar activity in the past. No extinction is found in excess of the galactic value. This points towards a currently quiet stellar environment and low activity in the central engine.
4.5 NGC 3783
Table 7: Parameters of the spectral synthesis for NGC 3783.
R (ring)
140-470 pc
$\sigma$ km/s 60
% stell 100$\pm$1
% PL 0
% BB -
T (BB) -
E(B-V) 0.01$\pm$0
F2V 33$\pm$1
M3V 42$\pm$1
F8III 5$\pm$0
M4III 18$\pm$2
M2I 2$\pm$1
Figure 11: HST image in the F160W filter of NGC 3783 with the ISAAC slit superimposed. The ring corresponds to the sum of the two regions labelled R.
Figure 12: Stellar population fit of the ring around the central region of NGC 3783 (see Fig. 11).
NGC 3783 is a nearly face-on SBa galaxy with an unobscured very bright and variable Seyfert 1 nucleus (e.g. Bentz et al., 2009). The sharp-divided image of NGC 3783 is shown in Fig. 11 with the studied region superimposed; it clearly shows the high contamination from the Sy1 nucleus in this galaxy.
Though NGC 3783 is extensively studied at all frequencies, the strong non-stellar luminosity in the centre makes difficult any quantitative study of the stellar population in the nucleus. Davies et al. (2007), from near-IR data, and Esquej et al. (2014), based on PAH measurements, could only provide an indication of the last burst around $\sim 70$ Myr. Based on the synthesis of a SOAR near-IR spectrum, Riffel et al. (2024) also inferred a hint of a young stellar population on top of the large contribution of an intermediate stellar population, leading to a mean age of $\sim$ 2 Gyr.
We made no attempt to recover the nuclear stellar population from our data in the central 140 pc, since the very strong AGN emission overwhelms the nuclear flux. As NGC 3783 is viewed almost face-on with a faint symmetrical bulge structure, a hereafter called ring-like spectrum is built by adding the very similar spectra extracted on either side of the nucleus (Boisson et al., 2002). The inferred stellar population is given in Table 7, and the corresponding spectrum and fit are shown in Fig. 12.
Though the stellar population is dominated by about 65% of old and evolved stars ($>$ 4.0 $\times$ 109yrs), a strong contribution of intermediate age stars (33%) is present as well as a hint of a recent star formation event. We may see the trace of a young population in this region surrounding the nucleus of NGC 3783.
4.6 NGC 6221
Table 8: Parameters of the spectral synthesis for NGC 6221.
C (nuc) N1 N2
65 pc 65-170 pc 170-355 pc
$\sigma$ km/s 60 70 70
% stell 42$\pm$1 100$\pm$0 100$\pm$0
% PL 58$\pm$1 0 0
% BB - - -
T (BB) - - -
E(B-V) - 0.65$\pm$0.05 0.97$\pm$0.04
A0V 1$\pm$0
F2V 5$\pm$1 17$\pm$1
rG1V 12$\pm$1 7$\pm$2
K4V 23$\pm$1
rM1V 1$\pm$0
wM1.5V 3$\pm$2
wM3V 11$\pm$2 20$\pm$1 17$\pm$2
F8III 18$\pm$1 3$\pm$0
K3III 2$\pm$2
K5-M0III 9$\pm$1
M4III 20$\pm$1 45$\pm$0 34$\pm$1
rK2I 7$\pm$1
rK4I 2$\pm$0
M2I 43$\pm$1
Figure 13: Sharp-divided IR HST image in the F160N filter of NGC 6221 with the ISAAC slit superimposed. The three regions are the nucleus (C) and the two regions to the north, N1 and N2.
Figure 14: Stellar population fits to the three regions of NGC 6221. From top to bottom: nucleus, N1 and N2 (see Fig. 13).
NGC 6221, is a barred Sbc galaxy located within a small galaxy group, with a bar clearly visible in the optical and IR. This X-ray bright Seyfert galaxy is known to exhibit clear signs of intense star formation together with a weak Seyfert nucleus, exhibiting both signatures of a Seyfert 1 in the X-rays and a typical reddened (Av=3) starburst-like spectrum in the optical (e.g. Levenson et al., 2001; La Franca et al., 2016).
The sharp-divided image of NGC 6221 is shown in Fig. 13 with the studied regions superimposed. We study the nuclear spectrum within 80 pc, and two regions towards the north up to 200 pc, excluding from our study the south conspicuous HII region spectrum (Boisson et al., 2002). The corresponding stellar populations are given in Table 8, with spectra and fits shown in Fig. 14. We note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
Although the stellar spectrum is strongly diluted (58% power-law contribution) in the very nucleus, one can see in Fig. 14 conspicuous stellar features due to the large contribution of M supergiants which have strong narrow features. We find 60% and 40% contributions of old stars and very young stars respectively. This young component confirms the presence of a very strong starburst on top of the AGN.
The stellar population at $\sim 100$ pc (N1) is old with hints of past star formation in a dusty environment suggested by the small intermediate age component. Further away, at $\sim 250$ pc from the nucleus (N2), in an even more dusty environment, it appears that a succession of bursts have occurred. Both regions exhibit a mixture of stars that are rich and poor in metals. We thus see a clear stellar population gradient together with a patchy extinction. All this is in coherence with the composite classification of NGC 6221.
4.7 NGC 7582
Table 9: Parameters of the spectral synthesis for NGC 7582.
S1 S2 N1 N2
40-120pc 120-215pc 40-120pc 120-215pc
$\sigma$ km/s 70 70 70 70
% stell 100$\pm$0 100$\pm$0 100$\pm$0 100$\pm$0
% PL 0 0 0 0
% BB - 2$\pm$0 - -
T (BB) - 200$\pm$20 - -
E(B-V) - - 0.52$\pm$0.17 0.78$\pm$0.02
A0V 2$\pm$1
F2V 58$\pm$1 44$\pm$2 25$\pm$1 35$\pm$1
F8III 20$\pm$1 11$\pm$1
K5-M0III 2$\pm$2 24$\pm$2 18$\pm$1 41$\pm$1
M4III 10$\pm$1 4$\pm$1 9$\pm$1
rF8Ib 1$\pm$1
rK4I 2$\pm$1
M2I 30$\pm$1 25$\pm$1 26$\pm$1 13$\pm$1
Figure 15: Sharp-divided IR HST image in the F606WN filter of NGC 7582 with the ISAAC slit superimposed. The north (N1, N2) and south (S1, S2) regions are indicated.
Figure 16: Stellar population fits to the four regions of NGC 7582. From top to bottom: N2, N1, S1, and S2 (see Fig. 15).
NGC 7582 is a nearby barred galaxy in the Grus quartet that harbours a nuclear starburst surrounding the active nucleus (Wold & Galliano, 2006; Bianchi et al., 2007). The Galactic column density towards NGC 7582 is very low. This X-ray bright Seyfert 2 galaxy (Bianchi et al., 2009) is a typical “changing look AGN”, a broad emission line component being reported by Aretxaga et al. (1999). This component was initially proposed as caused by a supernova explosion, but was later confirmed to be from an AGN (Ricci et al., 2018).
The sharp-divided image of NGC 7582 (Fig. 15) shows a disturbed morphology with an obscured nucleus, and a dust lane crossing over the nuclear region. Due to strong dilution by the AGN non-thermal continuum, the nuclear spectrum (80 arcsec diameter) is useless for stellar population studies. However, two external regions on each side of the nucleus were extracted, whose stellar populations are given in Table 9, with the corresponding spectra and fits shown in Fig. 16. The northern regions are shining through dust patches. This translates into higher E(B-V) in those regions than on the other side of the nucleus. We note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
We find an important population of evolved giant stars ($\sim(4-6)\times 10^{9}$ yrs), intermediate ones ($\sim 4\times 10^{8}$ yrs) as well as very young (107 yrs) stars in the four selected regions. This is in line with what is found in the optical range by Burtscher et al. (2021), and in the near IR by Riffel et al. (2024) within the 200 central parsecs.
In contrast with the traditional view that Seyfert 2s are essentially composed of an old stellar population, we do not find any cold dwarf (i.e. stars older than 1010 yrs) whatever the region sampled. Our results (neither dwarf nor metal-rich stars) points to a strong ongoing burst of star formation in the inner 200 pc. NGC 7582 may be at a star forming stage.
5 Summary and conclusions
Taking advantage of the rich near-IR H-band, we investigated the stellar populations of the central few 100  pc of the host galaxy of seven active galactic nuclei, with the help of a stellar spectral database that allowed us to detect the young star contribution down to $\leq 10^{7}$ yrs.
Our results can be summarised as follows:
First, in the inner 100 pc, we detect some metal-rich population in all AGNs but NGC 7582. The only Seyfert 1 for which we could synthesise in detail the nucleus, MCG 06-30-15, shows a strong nuclear starburst. The two composite Seyfert 2s, NGC 6221 and NGC 7582, are also dominated by a starburst component, while no young stars are detected in the genuine Seyfert 2 galaxies, NGC 2110, and NGC 3185, nor in NGC 2992.
Moreover, with our good spatial sampling, we could show how the star formation history of the inner nucleus of the galaxies is highly heterogeneous, as young bursts of star formation, intermediate (poststarburst) and old stellar populations all appear in widely varying proportions at all places. For example, one can note that in the four regions within 100 pc of MCG 06-30-15 (for which we have the best spatial and spectral resolutions), we find a large variety of populations. We could highlight such a diversity thanks to the IR domain, rich in stellar features, and where the star contribution is higher. Thus, there may not be only one process or relation between fuelling by star formation and power of the AGN.
Second, outside the nucleus ($>$100 pc) in both Seyfert 1 and Seyfert 2, we have resolved an intermediate or even young stellar population.
These results indicate that an active nucleus always seems to be linked to some past or even ongoing stellar formation. This was also noted by, for example, Davies et al. (2007) or Riffel et al. (2022) who, based on high spatial resolution integral field observations, found that the near-IR light in the inner few 100 pc of active galaxies is primarily composed on average of moderately young (a few $10^{7}$ to $10^{8}$ yrs) to intermediate-age (a few $10^{8}$ yrs) stellar populations, and that the age of the starburst seems connected to the AGN’s luminosity. However, those studies concern only the post-starburst population, with ages $\sim 10^{8}$$10^{9}$ yr.
Based on results of the stellar population study, in the optical, of a sample of 12 galaxies covering a range of morphology types and nuclear activities, Boisson et al. (2000) suggested the existence of a relationship between stellar population and activity type. The picture emerging from this study was that of an evolutionary sequence rather than a strict unified scheme for AGNs. Indeed, differences in the stellar populations within the nuclear regions were found to follow the degree of nuclear activity. Seyfert 2 galaxies appear to exhibit fossil star formation activity, that is, a stellar population younger than found in normal galaxies, but somewhat older and more metal-rich than found in starburst galaxies. In an optical study of the population synthesis of the inner 300 pc of star forming and active galaxies, including the three Seyfert 2 galaxies of our sample, Burtscher et al. (2021) found a dominance of young stellar populations too, while this was not the case for their control sample of inactive galaxies. They also showed that such young populations are not necessarily indicative of ongoing star formation. This is along the lines of the evolutionary sequence of Boisson et al. (2000, 2004). Based on the similar nuclear star formation histories of the Seyfert 2s and the star forming control sample galaxies, Burtscher et al. (2021) further speculate that the starbursts probably turned into Seyferts for some fraction of their time.
In the present paper, in the IR, we find that indeed star-forming galaxies may turn into Seyfert 1s, while star formation in Seyfert 2s, if any, appears outside the nuclear region. We note that NGC 2992, for example, is known to change from type 1 to type 2.
In conclusion, while there is little doubt that the above observations indicate that AGN activity and star formation do often coincide within the very centre of a given galaxy, we find a great diversity of stellar populations, and no clear correlation with the level of activity of the central engine or with the galaxy morphological type, most probably linked to the fact that several of these objects vary from Seyfert 1 to Seyfert 2 and even LINER.
Acknowledgements.
We thank the referee for comments and suggestions that helped us to improve the paper. I.M. acknowledges financial support from the State Agency for Research of the Spanish MCIU, through the “Center of Excellence Severo Ochoa” award to the Instituto de Astrofísica de Andalucía (CEX2021-001131-S), and through PID2022-140871NB-C21. F.D. acknowledges long-term support from CNES. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by1998 the National Aeronautics and Space Administration and operated by the California Institute of Technology, the NASA’s Astrophysics Data System bibliographic services, and of the Centre de Données de Strasbourg. We gratefully thank Sandra Raimundo who provided us with the VLT-SINFONI spectra of MCG-6-30-15.